OpenStudy (anonymous):

solve the porportion 7/y=3/17 round the answer to the nearest tenth

OpenStudy (anonymous):

\[y = \frac{119}{3}\approx4.0\]

OpenStudy (lalaly):

39.7

OpenStudy (anonymous):

7/y=3/17 --> 3y=119 --> y=119/3 (the ans is 39.7)

OpenStudy (anonymous):

3y =119 == > y =119/3=39.7

OpenStudy (anonymous):

39.7 sorry

OpenStudy (anonymous):

heyy its ok:)

OpenStudy (anonymous):

thnks